Solve the absolute value equation 7 l 2x-3 l =49
1
Expert's answer
2020-08-27T15:57:57-0400

7 l 2x-3 l =49


l 2x-3 l ="\frac{49}{7}"


 l 2x-3 l = 7


If |x| = a

Then x = a or x= -a


 l 2x-3 l = 7

2x - 3 = 7 or 2x - 3 = -7

2x= 7 +3 or 2x = -7 +3

2x= 10 or 2x=-4

x="\frac{10}{2}" or x="\frac{-4}{2}"

x=5 or -2

Thus, the solution of the given absolute value equation is x= 5 , -2.